The walls and ceiling inside your bedroom need to be painted. To save money, you decide that you will paint the bedroom yourself. Use the following information to solve this problem:
• The bedroom is 17 feet long by 18 feet wide, and the ceiling is 9 feet high.
• The inside of the bedroom door will be painted the same color as the walls.
• Two coats of paint will be applied to all of the painted surfaces.
• The room has one window, measuring 3 feet, 9 inches by 4 feet, which will not be painted.
Pólya’s Principle Step 1: Understand the Problem
1. Describe in detail what you understand the problem to be. In other words, what problem will you need to solve? Is there enough information to enable you to find a solution to your problem?

Answer by addingup(3677)

2 walls: 17 x 9 = 153 x 2 = 306 square feet
2 walls: 18 x 9 = 162 x 2 = 324 square feet
ceiling: 17 x 18 = 306
.
so we have a total area of 306 + 324 + 306 = 936 square feet
.
Now deduct the area of the window:
3 feet, 9 inches by 4 feet (convert to inches and calculate):
(36 + 9) x 48 = 2160 square inches. And convert to square feet:
2160/144 = 15 square feet
936 - 15 = 921 square feet is the area to be painted.
.
But you'll be painting it twice (2 coats) so 921 x 2 = 1842 square feet